# Why Transformers Need Adam: A Hessian Perspective

### v3

@@ 1 Introduction
2 Main Results
@@ 2.1 Problem Settings
Setups. Hessian of large-scale NNs are intractable to compute and store. In this work, we apply a numerical tool called Stochastic Lanczos Quadrature method (SLQ) [7] to approximate the Hessian spectrum. SQL uses a smooth curve on $\mathbb{R}$ to approximate the histograms of eigenvalues (see Figure LABEL:fig_full_spectrum as an example). A detailed introduction to SLQ is provided in Appendix C.2. All experimental setups in this section is shown in Appendix D. We focus primarily on the following models/tasks.
@@ 2.2 Full Hessian Spectrum Is Not Informative Enough
2.3 Main Findings Through Blockwise Hessian Spectra
What other factors could cause SGD to perform significantly worse than Adam on Transformers but not on CNNs? We identify some critical features that have been overlooked in the full Hessian spectrum analysis above.
  • 1.
    Hessian structure. Existing literatures showed that the Hessians of MLPs are close to block-diagonal matrices [18, 73, 60]. We restate their findings in Appendix B. Further, [18, Section 7] theoretically attributes the block diagonal structure to (i) the layer-by-layer design in NNs and (ii) the Cross-Entropy loss. Following this line of work, we also observe a near block-diagonal Hessian in small Transformers in Figure LABEL:fig_block_diagnal, where the variables in each principal block correspond to the parameters of each block in the Transformer. These results suggest that near block-diagonal Hessian might be common in NNs.
  • 2.
@@ 2.3 Main Findings Through Blockwise Hessian Spectra
    Build-up rules of Transformers. As shown in Figure LABEL:fig_blockwise_spectrum, CNNs are constructed by the repetitive stacking of similar parameter blocks (convolution layers). In contrast, Transformers consist of disparate parameter blocks, e.g. Query, Key, Value in attention, and MLP layers. Further, these blocks are stacked in a non-sequential manner. Due to the different designs of these blocks, they may have different properties for optimization, which can further affect the optimizer behavior.
Combining these together, we hypothesize that the blockwise Hessian spectrum, i.e., the spectrum of principal blocks of Hessian $[\nabla^{2}\mathcal{L}(w)]_{l},l\in[L]$, might provide additional insights. What extra information can be contained in the blockwise spectrum but not in the full spectrum? By definition, the blockwise Hessians form the principal block sub-matrix of the full Hessian. We note that Transformers are observed to have a near block-diagonal Hessian. For block diagonal matrices, blockwise spectra encode the location of eigenvalues, i.e., which block an eigenvalue (of the full matrix) resides in. In contrast, the full spectrum overlooks this location information. In the following, we study the blockwise Hessian spectra of various models and show that they indeed carry more information than the full spectrum for distinguishing CNNs and Transformers.
@@
We here demonstrate the shape of blockwise spectra in VGG16 [40] (CNN) and BERT [25] (Transformer). We sample four blocks for each model and present the spectra in Figure LABEL:fig_blockwise_spectrum. In BERT, the Hessian spectra of embedding, attention, and MLP blocks are largely different. In contrast, in ResNet, the spectra of convolution layers are similar. We further verify this observation for the rest of the parameter blocks. We calculate the Jensen-Shannon (JS) distance between two eigenvalue densities of all possible block pairs and show the results in Figure LABEL:fig_blockwise_heatmap.
@@
The results in Figure LABEL:fig_blockwise_heatmap showed a new phenomenon: for all Transformers we checked, the blockwise Hessian spectra are largely different across blocks. In the following, we refer to this phenomenon as “block heterogeneity". In contrast, the blockwise Hessian spectra of CNNs are similar and the block heterogeneity is not observed. We refer to this phenomenon as “block homogeneity". These results indicate that block heterogeneity is informative in distinguishing CNNs and Transformers. Intuitively, the block-homogeneity in CNNs comes from the repetitively similar convolution layers, while the block-heterogeneity in Transformers stems from the non-sequential stacking disparate layers such as Query, Value, and MLPs. In the following, we will show that the block heterogeneity is strongly correlated with the performance gap between SGD and Adam on Transformers.
@@ 2.4 SGD Performs Worse than Adam on Various Tasks with Block Heterogeneity
Figure LABEL:fig_blockwise_spectrum and LABEL:fig_blockwise_heatmap have shown that (1) SGD is worse than Adam on Transformers. (2) Transformers have block-heterogeneity. Now we further link block heterogeneity to SGD’s unsatisfactory performance on non-Transformer models. This would directly establish a connection between “block heterogeneity" and “why SGD is worse than Adam", without going through Transformers or attention blocks as an intermediary. We consider one man-made example and one real-world example.
@@
Example 2: MLP-mixer. We consider MLP-mixer [82], a famous all-MLP architecture that outperforms CNNs and ViTs on some vision tasks. Figure LABEL:fig_mlp_gap (b) (c) shows that the initial Hessian of MLP-mixer has block heterogeneity and SGD lags behind Adam on this architecture.
@@ 3.1 Experimental Observations
Summary of experimental observations. Figure LABEL:fig_quadratic presents two phenomena. For Hessian with heterogeneous blocks (Case 1 and 3), GD largely lags behind Adam. For Hessian with homogeneous blocks (Case 2 and 4), GD is on par with Adam. We emphasize that all Hessians have the same condition number. Further, Hessian of Case 3 and 4 share all the eigenvalues (not just the extreme ones). The gap between Adam and GD is due to the different blockwise spectra caused by the different locations of eigenvalues. We hypothesize that GD performs badly because it uses one single learning rate for all blocks, which cannot handle the heterogeneity among blocks. Such heterogeneity can be better handled using different learning rates across blocks, as designed in Adam.
@@ 3.2 Initial Theoretical Results
We now compare the complexity of Adam and that of GD. By Theorem 1, Adam is faster than GD when $r\cdot\max_{l\in[L]}\kappa_{l}\leq\kappa$. In the quadratic model with heterogeneous blocks (Case 3), our simulation over 1000 trials shows that $r\leq 1000$ with probability $\geq\frac{2}{3}$ when using standard Gaussian random initialization. Since $\max_{l\in[L]}\kappa_{l}\approx 1$, we have $r\cdot\max_{l\in[L]}\kappa_{l}\leq 1000$, w.h.p., and is about 5×5$\times$5 × smaller than $\kappa=5000$. So Adam could be 5×5$\times$5 × faster than GD, w.h.p.. This is indeed observed in Figure LABEL:fig_quadratic where Adam outperforms GD by a significant margin. We summarize the complexity of GD and Adam in Table 2. We remark that there is room to improve the complexity bound of Adam, but it could be challenging. We provide more technical discussions in Appendix B.
@@ 4 Conclusion
In this work, we explore why SGD largely underperforms Adam on Transformers. we establish a phenomenon called block heterogeneity in Hessian and link it to the performance gap between Adam and SGD. Initial theory is provided to support the claim.
@@ Appendix B More Results and Discussions
Block-diagonal structure in the existing literature.
We remark that Collobert [18], Roux et al. [73], Martens and Grosse [60] also observed the block-diagonal structure in (approximated) Hessian of small-scaled MLPs. We restate their findings in Figure LABEL:fig_block_diagnal_existing. [18, Section 7] further theoretically proved that the block diagonal structure stems from (i) the layer-by-layer structure and (ii) the Cross-Entropy loss. These results suggest that block-diagonal structure might be common in NNs.
@@
How to obtain a tighter complexity bound of Adam?
It is valid to ask whether the complexity upper bound in Theorem 1 $\kappa_{Adam,l}=r\kappa_{l}$ can be tightened, e.g., improve the factor of $r$. We point out it would be difficult if there is no extra structure on $H_{l}$. A key technical step is to bound the condition number of the preconditioned matrix $\kappa\left((D_{Adam,l}^{0})^{-1}H_{l}\right)$. Intuitively, a diagonal preconditioner of $H_{l}$ is powerful when $H_{l}$ itself has a near-diagonal structure, e.g., pure diagonal, tridiagonal or diagonal dominant [30]. Unfortunately, it is unclear whether these structures hold in Transformers. Without any assumption on $H_{l}$, we find that the diagonal preconditioner of $D_{Adam}^{0}$ could increase the condition number. For instance, when using standard Gaussian initialization, in case 3, we find $\kappa\left((D_{Adam,l}^{0})^{-1}H_{l}\right)$ equals $7.09\kappa_{1}$, $18.98\kappa_{2}$, $18.76\kappa_{3}$ for the 3 blocks, respectively (all averaged over 1000 trials). It would be interesting to explore if there are special structures of $H_{l}$ in Transformers such that Adam preconditioner can reduce $\kappa_{l}$, rather than increase it. We leave it as a future direction.
Although Adam preconditioner might not always reduce the “local" condition number $\kappa_{l}$, the coefficient in the complexity is now independent of the “global" condition number $\kappa$. As argued above, such changes in coefficient could lead to considerable improvement over GD. Such improvement in complexity is attributed to the block diagonal structure in Hessian as well as its heterogeneous blockwise spectrum. To our knowledge, such improvement is not shown in the existing literature. In summary, our theory indicates that: for problems with block heterogeneity, the single-learning rate methods like GD can largely lag behind coordinate-wise learning rate methods like Adam.

### v4

@@ 1 Introduction
2 Problem Settings and Initial Attempts
@@ 2.1 Problem Settings
Setups. Hessian of large-scale NNs are intractable to compute and store. In this work, we apply a numerical tool called Stochastic Lanczos Quadrature method (SLQ) [7] to approximate the Hessian spectrum. SQL uses a smooth curve on $\mathbb{R}$ to approximate the histograms of eigenvalues (see Figure LABEL:fig_full_spectrum as an example). A detailed introduction to SLQ is provided in Appendix C.2. All experimental setups are shown in Appendix D. We focus primarily on the following models/tasks.
@@ 2.2 Full Hessian Spectrum Is Not Informative Enough
2.3 Motivations of Investigating Blockwise Hessian Spectra
@@
What other factors could cause SGD to perform significantly worse than Adam on Transformers but not on CNNs? We identify one critical feature that has been overlooked in the full Hessian spectrum analysis above: the building-up rules of Transformers. As shown in Figure LABEL:fig_blockwise_spectrum, CNNs are constructed by the repetitive stacking of similar parameter blocks (convolution layers). In contrast, Transformers consist of disparate parameter blocks, e.g. Query, Key, Value in attention, and MLP layers. Further, these blocks are stacked in a non-sequential manner. We hypothesize that the “different designs among parameter blocks" can be reflected in the Hessian of these parameter blocks, which might affect algorithmic behavior. This inspires us to investigate the blockwise Hessian spectra, i.e., the spectrum of principal blocks of Hessian $[\nabla^{2}\mathcal{L}(w)]_{l},l\in[L]$.
In parallel to the motivation above, we further provide another evidence that blockwise spectra might be helpful. Classical literature showed that the Hessians of neural nets are near-block-diagonal matrices [18], i.e., the magnitudes in the Hessian principle blocks are much larger than those in the off-diagonal blocks. We restate their findings in Figure LABEL:fig_block_diagnal (a). This implies that the majority of Hessian information indeed lies in its principle blocks, and the blockwise Hessian of neural nets might contain valuable information.
To summarize, the “heterogeneous" building-up rules of Transformers inspire us to check the blockwise Hessian, i.e., the principle blocks of the Hessian. The classical results of neural nets [18] further support us to explore this direction since they find that the majority of Hessian information indeed lies in its principle blocks. In the following, we study the blockwise Hessian spectra of various neural networks. For ease of implementation, we define parameter blocks under the PyTorch partition. We show that the blockwise spectra indeed carry more information than the full spectrum for distinguishing CNNs and Transformers.
Remark: why near-block-diagonal? We briefly restate the analysis in [18, Section 7] to explain the near-block-diagonal Hessian structure of neural nets. Consider minimizing $\ell(f(\theta,x),y)$ where $\ell(\cdot,\cdot)$ is the Cross-Entropy (CE) loss, $f(\theta,x)=\sum_{i=1}^{n}v_{i}\phi(w_{i}^{\top}x)$ is an 1-hidden-layer neural network with input $x\in\mathbb{R}^{d}$, weight $w_{i}\in\mathbb{R}^{d}$, $v_{i}\in\mathbb{R}$, and label $y\in\{0,1\}$, then the off-diagonal-block Hessian elements will contain
$\frac{\partial^{2}\ell(f(\theta,x),y)}{\partial w_{i}\partial w_{j}}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}p_{\theta}(y|x)\left(1-p_{\theta}(y|x)\right)}v_{i}v_{j}\phi^{\prime}\left(w_{i}^{\top}x\right)\phi^{\prime}\left(w_{j}^{\top}x\right)xx^{\top}\quad\text{for }i\neq j,$ (1)
where $p_{\theta}(y|x)=1/(1+\exp(-yf(\theta,x)))$ and $\phi^{\prime}(\cdot)$ is the derivative of $\phi(\cdot)$. Note that the term $p_{\theta}(y|x)\left(1-p_{\theta}(y|x)\right)$ will vanish rapidly since the training objective is to maximize $p_{\theta}(y|x)$. Consequently, this drives the Hessian towards a near-block-diagonal configuration, with each block representing an output neuron. This result is validated in Figure LABEL:fig_block_diagnal: we find that the near-block-diagonal structure appears at 1% step and it maintains along the training.
3 Main Results
3.1 Transformers Exhibit Block Heterogeneity in Hessian, while CNNs Do Not
@@
We now compare the shape of blockwise spectra in VGG16 [40] (CNN) and BERT [25] (Transformer). We sample four blocks for each model and present the spectra in Figure LABEL:fig_blockwise_spectrum. In BERT, the spectra of embedding, attention, and MLP blocks are largely different. In contrast, in ResNet, the spectra of convolution layers are similar. We further verify this observation for the rest of the parameter blocks. We calculate the Jensen-Shannon (JS) distance between two eigenvalue densities of all possible block pairs and show the results in Figure LABEL:fig_blockwise_heatmap. We summarize our findings in Observation 1.
Observation 1: For all Transformers we checked, the blockwise Hessian spectra are largely different from each other. In contrast, the blockwise Hessian spectra of CNNs are similar.
@@
In the following, we refer to the phenomenon of Transformers as “block heterogeneity", and refer to that of CNN as “block homogeneity". The observations in Figure LABEL:fig_blockwise_spectrum and LABEL:fig_blockwise_heatmap indicate that block heterogeneity is informative in distinguishing CNNs and Transformers. In the following, we will show that the block heterogeneity is strongly correlated with the performance gap between SGD and Adam on Transformers.
@@ 3.2 SGD Performs Worse than Adam on Various Tasks with Block Heterogeneity
Figure LABEL:fig_blockwise_spectrum and LABEL:fig_blockwise_heatmap have shown that (1) SGD is worse than Adam on Transformers. (2) Transformers have block heterogeneity. Now we further link block heterogeneity to SGD’s unsatisfactory performance on non-Transformer models. This would directly establish a connection between “block heterogeneity" and “why SGD is worse than Adam", without going through Transformers or attention blocks as an intermediary. We consider one man-made example and one real-world example.
@@
Example 2: MLP-mixer. We consider MLP-mixer [81], a famous all-MLP architecture that outperforms CNNs and ViTs on some vision tasks. Figure LABEL:fig_mlp_gap (b) (c) show that the initial Hessian of MLP-mixer has block heterogeneity and SGD lags behind Adam on this architecture.
We summarize the findings so far in Observation 2.
Observation 2: For all tasks that we checked, SGD is worse than Adam when block heterogeneity exists, regardless of whether Transformers or attention mechanisms are utilized.
@@ 3.3 Reduced Block Heterogeneity in Pre-trained Transformers
In Figure LABEL:fig:hessian_along_training_vit in Appendix B, we further report the evolution of the block heterogeneity of ViT-base along the training. Similarly to GPT2 in Figure LABEL:fig_sft_gap, we find that the block heterogeneity of ViT-base tends to reduce after the training. In addition, we find that SGD can perform better when initializing at the weight with less heterogeneity, e.g., initializing at 50% total training steps. We hypothesize that “the attenuation of Hessian heterogeneity" is a common phenomenon after training, and we leave detailed investigation as a future direction.
Observation 3: Block heterogeneity in Hessian tends to reduce after (pre)-training.
@@ 3.4 Implication on Choosing SGD or Adam
Finally, we remark $JS^{0}$ is rather expensive to compute due to the overhead of SQL: it requires comparable time to one training run. Fortunately, we find the original SQL is rather redundant for measuring hessian heterogeneity. We propose some simple tricks to significantly reduce the computation time, while still effectively detecting the Hessian heterogeneity. We call it simplified SQL and we present it in Table 3 in Appendix B. As a result, the simplified SQL can obtain the same message as in Table 1 while only taking negligible time (e.g., $<0.001$s for ResNet18).
@@ 4.1 Experimental Observations
Summary of experimental observations. Figure LABEL:fig_quadratic presents two phenomena. For Hessian with heterogeneous blocks (Case 1 and 3), GD largely lags behind Adam. For Hessian with homogeneous blocks (Case 2 and 4), GD is on par with Adam. We emphasize that all Hessians have the same condition number. Further, Hessian in Case 3 and 4 share all the eigenvalues (not just the extreme ones). The gap between Adam and GD is purely due to the different blockwise spectra caused by the different locations of eigenvalues. Case 3 and 4 help reveal the causal relation between “block heterogeneity in Hessian" and “GD is worse than Adam". We hypothesize that GD performs badly because it uses one single learning rate for all blocks, which cannot handle the heterogeneity among blocks. Such heterogeneity can be better handled using different learning rates across blocks, as designed in Adam.
@@ 4.2 Initial Theoretical Results
We now compare the complexity of Adam and that of GD. By Theorem 1, Adam is faster than GD when $r\cdot\max_{l\in[L]}\kappa_{l}\leq\kappa$. In the quadratic model with heterogeneous blocks (Case 3), our simulation over 1000 trials shows that $r\leq 1000$ with probability $\geq\frac{2}{3}$ when using standard Gaussian random initialization. Since $\max_{l\in[L]}\kappa_{l}\approx 1$, we have $r\cdot\max_{l\in[L]}\kappa_{l}\leq 1000$, w.h.p., and is about 5×5$\times$5 × smaller than $\kappa=5000$. So Adam could be 5×5$\times$5 × faster than GD, w.h.p.. This is indeed observed in Figure LABEL:fig_quadratic where Adam outperforms GD by a significant margin. We summarize the complexity of GD and Adam in Table 2.
Remark: some common misconceptions. During the review process, we find that readers might conclude that “Theorem 1 implies Adam under homogeneity has worse complexity than Adam under heterogeneity". We now clarify that this claim is not correct, and there is no conclusion on “whether Adam under homogeneity is faster or slower than Adam under heterogeneity". Similarly, Theorem 1 does not imply “Adam always converges similarly as GD under homogeneity". Though it is observed on CNNs, there is no general conclusion of this kind. For interested readers, we provide a detailed explanation in Appendix B.
@@
How to obtain a tighter complexity bound of Adam?
It is valid to ask whether the complexity upper bound in Theorem 1 can be tightened, e.g., improve the factor of $r$. We point out it would be difficult if there is no extra structure on $H_{l}$. A key technical step is to bound the condition number of the preconditioned matrix $\kappa\left((D_{Adam,l}^{0})^{-1}H_{l}\right)$. Intuitively, a diagonal preconditioner of $H_{l}$ is powerful when $H_{l}$ itself has a near-diagonal structure, e.g., pure diagonal, tridiagonal or diagonal dominant [30]. Unfortunately, it is unclear whether these structures hold in Transformers. Without any assumption on $H_{l}$, we find that the diagonal preconditioner of $D_{Adam}^{0}$ could increase the condition number. For instance, when using standard Gaussian initialization, in case 3, we find $\kappa\left((D_{Adam,l}^{0})^{-1}H_{l}\right)$ equals $7.09\kappa_{1}$, $18.98\kappa_{2}$, $18.76\kappa_{3}$ for the 3 blocks, respectively (all averaged over 1000 trials). It would be interesting to explore if there are special structures of $H_{l}$ in Transformers such that Adam preconditioner can reduce $\kappa_{l}$, rather than increase it. We leave it as a future direction.
More discussions on the theoretical advantage of Adam. Although Adam preconditioner might not always reduce the “local" condition number $\kappa_{l}$, the coefficient in the complexity is now independent of the “global" condition number $\kappa$. As argued above, such changes in coefficient could lead to considerable improvement over GD. Such improvement in complexity is attributed to the block diagonal structure in Hessian as well as its heterogeneous blockwise spectrum. To our knowledge, such improvement is not shown in the existing literature. One possible reason is that: for the optimization community, it is very rare to analyze (near-) block-diagonal Hessian structure since typical problems do not have such structure. For instance, in the classical non-linear programming dataset [48], all problems have non-block-diagonal Hessian. We suggest a different perspective to characterize modern optimization problems. We believe our perspective is new because it is built upon multiple non-trivial findings.
In summary, our theory indicates that: for problems with block heterogeneity, the single-learning rate methods like GD can largely lag behind coordinate-wise learning rate methods like Adam.
@@ 5 Conclusion
In this work, we explore why SGD largely underperforms Adam on Transformers. we establish a phenomenon called block heterogeneity in Hessian and link it to the performance gap between Adam and SGD. We numerically verify our claim on various Transformers, CNNs, MLPs, and quadratic problems. Initial theory is also provided to support the claim.
Acknowledgements
Yushun Zhang would like to thank Yinyu Ye, Wentao Ding, Guiyu Hong, Yingru Li, and Bohan Wang for the valuable discussions. The work of Ruoyu Sun is supported by NSFC (No. 12326608); Hetao Shenzhen-Hong Kong Science and Technology Innovation Cooperation Zone Project (No.HZQSWS-KCCYB-2024016); University Development Fund UDF01001491, the Chinese University of Hong Kong, Shenzhen; Guangdong Provincial Key Laboratory of Mathematical Foundations for Artificial Intelligence (2023B1212010001). The work of Z.-Q. Luo was supported by the Guangdong Major Project of Basic and Applied Basic Research (No.2023B0303000001), the Guangdong Provincial Key Laboratory of Big Data Computing, and the National Key Research and Development Project under grant 2022YFA1003900.
Broader Impacts
We explore why SGD performs worse than Adam for training Transformers. Our work can help the community better understand large AI model training. However, it would be a potential threat if the AI models are used for illegal usage.
@@ Appendix B More Results and Discussions
More results for SGD under careful tuning.
For ViT-base, GPT2-nano, and BERT, we further present the performance of SGD under learning rate grid search. For ViT-base training on ImageNet, we report the results after 30 epochs (or equivalently, about 30k iterations). We cannot afford further training ViT-base due to the limited hardware resources (a complete run of 90 epochs would take $>2$ weeks for each curve). As shown in Figure LABEL:fig:sgd_grid_search, SGD consistently performs worse than Adam on ViT-base, GPT2, and BERT. The best results for SGD are picked out and presented in Figure LABEL:fig:nlp_figure.
Note that we are not the first ones to report that “SGD performs worse than Adam on ViT". An influential work [92] also reports that SGD is worse than Adam on vanilla ViT. The authors report that “SGD yields significantly worse results than AdamW (on ViT)", and “ViT often fails to converge with SGD " (their Figure 3). These results align with our findings in Figure LABEL:fig:sgd_grid_search.
Detailed training curves of Figure LABEL:fig_mlp_gap.
In Figure LABEL:fig_mlp_gap_appendix, we present the detailed training curves for the man-made MLP in Figure LABEL:fig_mlp_gap. We find that SGD performs worse as heterogeneity grows, while Adam still performs well.
The evolution of Hessian heterogeneity along training.
For ViT-base, we further investigate the evolution of block heterogeneity of Hessian along the training. As shown in Figure LABEL:fig:hessian_along_training_vit, we find that heterogeneity attenuates along the training. We further take the checkpoint of ViT-base at 50% training step and switch AdamW to SGD, we observe that now SGD performs better than training from scratch as in Figure LABEL:fig:nlp_figure (a). SGD performs better here because there is less heterogeneity when initializing in the middle of training.
Simplifed SQL for calculating $JS^{0}$ in Section 3.4.
We note that $JS^{0}$ in Table 1 is rather expensive to compute due to the computational overhead of SQL: it requires comparable time to one training run. Fortunately, we find the original SQL is redundant for measuring hessian heterogeneity. We propose the following simple tweaks to significantly reduce the computation time, while still effectively detecting the Hessian heterogeneity. We call it simplified SQL.
  • Change the hyperparemters of SQL, including:
    • We change $\texttt{num}_{v}=10$ to $\texttt{num}_{v}=1$. In SQL, $\texttt{num}_{v}$ decides the number of random Gaussian vectors to approximate the expected quadrature. It is reasonable to reduce $\texttt{num}_{v}$ because in high dimensional space, random vectors tend to concentrate around their mean, so one random sample can already be informative enough.
    • We change the Lanzcos step $m=100$ to $m=10$. The reduction on Lanzcos step will have a coarse estimation on the middle eigenvalue, but won’t affect much the heterogeneity, which is more dependent on the extreme eigenvalues.
  • Randomly sample a subset of blocks and reduce batch size for estimating the spectrum. We uniformly sample 50% blocks and choose batch size = 32 (previously batch size = 1024).
We report the result and runtime in Table 3. As a result, the simplified SQL can obtain the same message as the original SQL: $JS^{0}$ of ResNet is about 100x smaller than that of BERT. Further, the simplified SQL is highly efficient to compute. With this simplified SQL, we believe our method can efficiently scale to larger models. The result is tested on a single V100.
Table 3: $JS^{0}$ computed by simplified SQL. We find that the simplified SQL can obtain the same message as the original SQL: $JS^{0}$ of ResNet is about 100x smaller than that of BERT. Further, the simplified SQL is efficient to compute.
Model JSO Time for JSO Time for Training Time for JSO / Time for training
BERT 98.8344 20 s 4 h 0.0014
ResNet18 0.3569 65 s 87.5 h 0.0002
@@
Does our theory imply that Adam under homogeneity is slower than Adam under heterogeneity?
 [^5] We would like to thank the anonymous Reviewer pTmy for raising this insightful question.
During the review process, we find that readers might conclude that “Theorem 1 implies Adam under homogeneity has worse complexity than Adam under heterogeneity". We would like to clarify that this conclusion is not correct, and there is no conclusion on “whether Adam under homogeneity is faster or slower than Adam under heterogeneity ". We explain as follows.
Our theoretical result states that Adam has complexity $\mathcal{O}\left(\max_{l}\kappa_{l}\right)$. If our result implies the above conclusion, one needs the following argument: when changing heterogeneity to homogeneity, $\max_{l}\kappa_{l}$ increases, and thus Adam is slower. However, “changing heterogeneity to homogeneity" does not necessarily mean “$\max_{l}\kappa_{l}$ increases". Actually, $\max_{l}\kappa_{l}$ can change in an arbitrary way (can increase, decrease, or keep the same) when changing the heterogeneity. We provide three examples below.
We will use Adam (homo) to denote the convergence rate of Adam on homogeneous Hessian, similarly for Adam (hetero).
Example 1: Adam (homo) is same as Adam (hetero).
  • Case 1-1 (homogeneous): eigenvalues are {1,2}, {1,2}
  • Case 1-2 (heterogeneous): eigenvalues are {1,2}, {11,12}
Since $\max_{l}\kappa_{l}$ are the same for both Case 1-1 and 1-2, Adam (homo) is the same as Adam (hetero).
Example 2: Adam (homo) is faster than Adam (hetero).
  • Case 2-1 (homogeneous): eigenvalues are {1,1.5}, {1,1.5}
  • Case 2-2 (heterogeneous): eigenvalues are {1,2}, {11,12}
Since Case 2-1 has smaller $\max_{l}\kappa_{l}$ than Case 2-2, Adam (homo) is faster than Adam (hetero).
Example 3: Adam (homo) is slower than Adam (hetero).
  • Case 3-1 (homogeneous): eigenvalues are {1,12}, {1,12}
  • Case 3-2 (heterogeneous): eigenvalues are {1,2}, {11,12}
Since Case 3-1 has larger $\max_{l}\kappa_{l}$ than Case 3-2, Adam (homo) is slower than Adam (hetero).
To sum up, there is no conclusion on “whether Adam under homogeneity is faster or slower than Adam under heterogeneity ". Either case can happen.
One possible source of confusion may come from the numerical examples (Case 3 and 4 in Section Section 4.1). If comparing two figures, Adam (homo) in Case 3 is slower than Adam (hetero) in Case 4. But as argued above, this is just one example, and it does not show Adam (homo) is always slower than Adam (hetero).